OpenStudy (anonymous):

What is the difference between negative and positive? Isn't 0 positive?

OpenStudy (compassionate):

Zero IS an integer, but it is neither positive or negative. It is sometimes called "non-negative," but it is not positive. Positive integers are also called "natural" or "counting" numbers. You use them to count: 0,1,2,3,4,5... Negative integers are the "opposite" of counting numbers. The opposite of 1 is minus(negative) 1. The opposite of 2 is minus 2. And so on. These negative numbers are steps backwards on a number line. Here are more opinions and answers from other FAQ Farmers: An integer is any whole number that doesn't contain anything beyond (to the right of) the decimal point. Positive and Negative are adjectives that describe the integer. Positive means greater than or equal to 0 (eg. 0,1,2,3,4,5...) and negative means less than or equal to 0 (eg. 0,-1,-2,-3,-4,-5...). Negative integers are preceded with a "-" sign except for 0 because -0=0. The set of positive integers and the set of negative integers when used by themselves act the same way under addition (they are just spelt a little differently) 3+3=6 -3-3=(-3)+(-3)=(-6) They act differently under multiplication though. 4*5=20 (-4)*(-5)=20 (not -20) but (-4)*5=4*(-5)=-20 :CITE: Read more: http://wiki.answers.com/Q/What_is_the_difference_between_positive_integers_and_negative_integers#ixzz21QexxPmd

OpenStudy (anonymous):

0 is neither negative nor positive.. and also 0 is neither even nor odd..

OpenStudy (zzr0ck3r):

0 is even 2(0) = 0

OpenStudy (zzr0ck3r):

every even number can be written as 2k where k is some integer.

OpenStudy (compassionate):

Stop it!

OpenStudy (zzr0ck3r):

?

OpenStudy (compassionate):

Sorry, zzr, that wasn't directed at you. You know who you are.

OpenStudy (zzr0ck3r):

every integer is even or odd

OpenStudy (compassionate):

Zero isn't an integer.

OpenStudy (zzr0ck3r):

yes it is

OpenStudy (anonymous):

Yes it is..

OpenStudy (compassionate):

0 has no value. It can't possibly be an integer.

OpenStudy (zzr0ck3r):

and its even, you guys go look it up:)

OpenStudy (callisto):

@Compassionate Your last saying contradicts to your first comment here... ''Zero IS an integer'' and ''Zero isn't an integer. ''

OpenStudy (zzr0ck3r):

:)

OpenStudy (compassionate):

I say it's more of an argument.

OpenStudy (zzr0ck3r):

no its not, its defined

OpenStudy (compassionate):

I'm more concerned on how can you have an integer when it has no value?

OpenStudy (zzr0ck3r):

how can you have a number when it has no value? 0 is not a number

OpenStudy (zzr0ck3r):

Im pretty sure at this point your messing with us

OpenStudy (compassionate):

Pardon me, I got my definition messed up. You're right, Zzr.

OpenStudy (compassionate):

Well this thread died.

OpenStudy (anonymous):

isn't a 0 number??

OpenStudy (unklerhaukus):